Find an equation of the line L that passes through the point
(−2, 4)
and satisfies the given condition. (Let x be the independent variable and y be the dependent variable.)
L is perpendicular to the line
2x + 5y = 10

Answer by flame8855(424)

L : 5y = -2x + 10
y = -2/5x +2
slope = -2/5
aa' =-1
a' = 5/2
y-4/x+2 = 5/2
y-4 = 5/2x +5
y = 5/2x +9